OpenStudy (anonymous):

1. Write 41n 6-21n 2 as a single natural logarithm. (1 point)

OpenStudy (zehanz):

Use 2 rules: \[a \ln b=\ln b^a\]\[\ln a - \ln b = \ln \frac{a}{b}\]

OpenStudy (anonymous):

that dont help i dont understand thoes rules

OpenStudy (zehanz):

First rule means: if you have 4 ln 6, you can also write it as:\[\ln 6^4\] Also: 2 ln 2 is the same as:\[\ln 2^2 = \ln 4\] So your problem now is a little smaller:\[\ln 6^4 - \ln 4\] The seond rule says: \[\ln 6^4 - \ln 4 = \ln \frac{ 6^4 }{ 4 }\]If you work out what \[\frac{ 6^4 }{ 4 }\]is, you have the answer.

OpenStudy (anonymous):

ok ty

OpenStudy (zehanz):

YW!